A cart is travelling over an xy plane with acceleration components ax = 4.0 m/s^2 and ay=-2.0 m/s^2. if the initial velocity along x is 8.0 m/s and along y is 12 m/s. find, the displacement after 3.5 seconds


1
Expert's answer
2021-09-20T16:31:57-0400

The displacement along x axis is given as follows:


"d_x = v_{x0}t+\dfrac{a_xt^2}{2}"

where "v_{x0} = 8m\/s" is the initial velocity along the x-axis, and "a_x = 4m\/s^2" is the acceleration along the x-axis. Thus, for "t = 3.5s", obtain:


"d_x = 8m/s\cdot 3.5s+\dfrac{4m/s^2\cdot (3.5s)^2}{2} \approx 53m"

Similarly for y-axis:


"d_y = v_{y0}t+\dfrac{a_yt^2}{2}\\
d_y = 12m/s\cdot 3.5s-\dfrac{2m/s^2\cdot (3.5s)^2}{2}\approx 30m"

Thus, the displacement vector is:


"\mathbf{d} = (d_x, d_y) = (53, 30)\space m"

Answer. "\mathbf{d} =(53, 30)\space m".